Project A is currently being considered by your company. It has the following projected cash flows:

                Year                       Project A                              

                  0                         -$300,000            

                  1                              90,000                                       

                  2                              90,000                                     

                  3                             110,000                                       

                  4                             110,000

The required rate of return for this project is 10 percent.

Document the FV you come up with in Step #1.  

Document the following for Step #2:

FV

PV

N

PMT

I/Y

Accept or Reject?
Answer #1

Initial Project Investment = I = $ 300000, Annual Project Cash Flows: Year 1 = $ 90000, Year 2 = $ 90000, Year 3 = $ 110000 and Year 4 = $ 110000

Required Rate of Return = 10 %

Future Value = FV = Total Present Value of Annual Project Cash Flows compounded at the required rate of return of 10%.

Present Value = Total Present Value of Annual Project Cash Flows discounted at the required rate of return of 10%.

N = Project Tenure = 4 years

PMT = Annual Project Cash Flows, PMT for Year 1 and Year 2 = $ 90000 and PMT for Year 3 and Year 4 = $ 110000

I/Y = Required Rate of Return = 10 %

The decision to accept or reject the project will depend upon the net present value (NPV) of the project which is calculated as shown below:

NPV = PV - Initial Investment

PV = 90000 / (1.1) + 90000 / (1.1)^(2) + 110000 / (1.1)^(3) + 110000 / (1.1)^(4) = $ 313974.455

Initial Investment = $ 300000

NPV = 313974.455 - 300000 = $ 13974.455

As the project has a positive NPV, the project should be accepted.